# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance

### v1

@@ IX-E Proof of Lemma 4
When the state $\xi$ enters in the jump set of the move-to-target mode $\mathcal{J}_{0}$ (20), the control law switches to the obstacle-avoidance mode by updating the mode indicator variable $m(t_{1},j_{1}+1)$ to $1$, using (27). Hence, for all time $(t,j)\in({I_{j_{1}+1}}\times j_{1}+1)$, the control vector $\mathbf{u}(\xi(t,j))$ in (11a) is given by
$\mathbf{u}(t,j)=\kappa_{r}\mathbf{v}(\mathbf{x}(t,j),\mathbf{a}(t,j)),$ (43)
where $\mathbf{a}(t,j)=\mathbf{a}(t_{1},j_{1}+1)$ as per (27). Observe that for all $(t,j)\in({I_{j_{1}+1}}\times j_{1}+1)$, one has $\mathbf{v}(\mathbf{x}(t,j),\mathbf{a}(t,j))\in\mathcal{P}(\mathbf{x}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$ and, according to (27), $\mathbf{h}(t_{1},j_{1}+1)=$ $\mathbf{x}(t_{1},j_{1}+1)$. Hence, for all $(t,j)\in({I_{j_{1}+1}}\times j_{1}+1)$,
$\mathbf{x}(t,j)\in\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1)).$ (44)
Since $\mathbf{x}(t_{1},j_{1})\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma_{s}}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ and $\mathbf{a}(t_{1},j_{1}+1)\in\mathcal{P}^{\perp}(\mathbf{x}(t_{1},j_{1}))$, the hyperplane $\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$ has a non-empty intersection with the interior of the $r_{a}-$dilated obstacle $\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$ i.e., $\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\neq\emptyset.$ As a result, it can be shown that there does not exist a point $\mathbf{q}$ in the set $\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))\setminus\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})$ such that the vector $\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))(\mathbf{q}-\Pi(\mathbf{q},\mathcal{O}_{i}))=\mathbf{0}.$ Moreover, for any non-zero vectors $\mathbf{z}\in\mathbb{R}^{3}\setminus\{\mathbf{0}\}$ and $\mathbf{s}\in\mathbb{R}^{3}\setminus\{\mathbf{0}\}$, the inner product $\mathbf{z}^{\intercal}\mathbf{P}(\mathbf{s})\mathbf{z}$ is always non-negative and is zero only if $\mathbf{P}(\mathbf{a})\mathbf{z}=0.$ Since $\mathbf{x}(t,j)\in\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$ and $\mathbf{a}(t,j)=\mathbf{a}(t_{1},j_{1}+1)$ for all $(t,j)\in({I_{j_{1}+1}}\times j_{1}+1)$, one has $\mathbf{x}_{\pi}(t,j)^{\intercal}\mathbf{P}(\mathbf{a}(t,j))\mathbf{x}_{\pi}(t,j)>0$, where $\mathbf{x}_{\pi}(t,j)=\mathbf{x}(t,j)-\Pi(\mathbf{x}(t,j),\mathcal{O}_{\mathcal{W}}),$ with $\Pi(\mathbf{x}(t,j),\mathcal{O}_{\mathcal{W}})=\Pi(\mathbf{x}(t,j),\mathcal{O}_{i}),$ and claim 1 in Lemma 4 is proved.
Next, we shows that when the control input corresponds to the obstacle-avoidance mode,
@@
for all $\mathbf{x}\in\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))\cap\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. This, combined with the facts that the control vector trajectory $\mathbf{u}(\xi)$ is continuous, when it corresponds to the obstacle-avoidance mode, according to Lemma 2, ensures that $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
@@
Note that $\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cup\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i}).$ We know that for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\subseteq\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$, where $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ with $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})=\Pi(\mathbf{x},\mathcal{O}_{i}).$ Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$, the control vector (11a) is given by $\mathbf{u}(\xi)=\kappa_{r}{\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi}}$. Moreover, according to claim 1 in Lemma 4, one has $\mathbf{x}_{\pi}^{\intercal}\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi}>0$. Hence, the control vector $\mathbf{u}(\xi)$ belongs to the open half-space $\mathcal{P}_{>}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$. As a result, in this case, for all $\mathbf{x}\in\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))\cap\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, condition (45) holds true.
@@
It is clear that for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})\subseteq\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$, the control vector (11a) is given by $\mathbf{u}(\xi)=-\kappa_{r}{\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi}}$ Moreover, according to claim 1 in Lemma 4, it is clear that $\mathbf{x}_{\pi}^{\intercal}\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi}>0$. Hence, the control vector $\mathbf{u}(\xi)$ belongs to the open half-space $\mathcal{P}_{<}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$. As a result, in this case, for all $\mathbf{x}\in\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))\cap\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, condition (45) holds true. Hence, using (44) and (45), claim 2 in Lemma 4 is proved.
@@
Next, we proceed to prove claim 3 in Lemma 4 which implies that when $\xi(t_{1},j_{1}+1)\in\mathcal{F}_{1}$ for some $(t_{1},j_{1}+1)\in\text{ dom }\xi$, the control input steers the state $\xi$ to the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}$ in finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty.$
@@
Let us define the set $\mathcal{O}_{i}^{\mathcal{S}}=\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$, as shown in Fig. 5. Since obstacle $\mathcal{O}_{i}$ is convex, the set $\mathcal{O}_{i}^{\mathcal{S}}$ is also convex. As a result, the target location has a unique closest point on the set $\mathcal{O}_{i}^{\mathcal{S}}$, represented by $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}}).$ Let us define the set $\mathcal{LS}:=\mathcal{L}_{s}(\mathbf{0},\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}}))\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. Since $\mathbf{0}\in\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1)),$ the line segment $\mathcal{LS}$ belongs to the hyperplane $\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1)).$ Note that the line segment $\mathcal{LS}$ also belongs to the exit region $\mathcal{R}_{e}$. Since the hit point $\mathbf{h}(t_{1},j_{1}+1)$ belongs to $\mathcal{R}_{l}^{i}$, the target location $\mathbf{0}$ is closer to the location $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}})$ than to the hit point $\mathbf{h}(t_{1},j_{1}+1)$. Hence, for a sufficiently small value of $\bar{\epsilon}$, used in (22), one can ensure that the set $\mathcal{LS}$ belongs to the set $\mathcal{J}_{1}^{\mathcal{W}}$.
@@
Now, if one ensures that the state $\mathbf{x}$, which belongs to the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$ after time $(t_{1},j_{1})$, in the obstacle-avoidance mode around obstacle $\mathcal{O}_{i}$, eventually intersects the set $\mathcal{LS}$ at some finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$, then it will imply that $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$, and claim 3 in Lemma 4 will be proven. To that end, let us divide the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$, as shown in Fig. 5, into 3 separate subsets as follows:
@@
where $\mathbf{h}=\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}=\mathbf{a}(t_{1},j_{1}+1)$, and the sets $\mathcal{S}_{1},\mathcal{S}_{2}$ and $\mathcal{S}_{3}$ are defined as follows:
@@
Figure 5: The partition of the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$.
@@
We show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ belongs either in the set $\mathcal{S}_{1}$ or in the set $\mathcal{S}_{3}$, the control eventually steers the state $\mathbf{x}$ to the set $\mathcal{S}_{2}.$ Then, we show that for all $\mathbf{x}\in\mathcal{S}_{2}$, the control vector $\mathbf{u}(\xi)$ belongs to the open positive half-space $\mathcal{P}_{>}(\mathbf{0},\mathbf{R}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi})$. This implies that the state $\mathbf{x}$, which belongs to the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h}(t_{1},j_{1}+1),\mathbf{a}(t_{1},j_{1}+1))$ after time $(t_{1},j_{1})$, in the obstacle-avoidance mode around obstacle $\mathcal{O}_{i}$, is always steered to the open positive half-space $\mathcal{P}_{>}(\mathbf{x},\mathbf{R}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi})$ and will eventually reach the set $\mathcal{LS}$ at some finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1).$
@@
First we show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ is either in the set $\mathcal{S}_{1}$ or in the set $\mathcal{S}_{3}$, the control will eventually steer the state $\mathbf{x}$ to the set $\mathcal{S}_{2}.$ When the control input corresponds to the obstacle-avoidance mode and $\mathbf{x}$ belongs to the set $\mathcal{S}_{1}$, the control vector $\mathbf{u}(\xi)$ in (11a) becomes
$\mathbf{u}(\xi)=\kappa_{r}{\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi}},\kappa_{r}>0.$ (48)

### v2

@@ IX-E Proof of Lemma 4
First, we prove that when the control input corresponds to the obstacle-avoidance mode, $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$. To that end, we make use of Nagumo’s theorem [8, Theorem 4.7] and show that when the control input corresponds to the obstacle-avoidance mode, one has
@@
for all $\mathbf{x}\in\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. This, combined with the fact that the control vector trajectory $\mathbf{u}(\xi)$ is continuous, when it corresponds to the obstacle-avoidance mode, as stated in Lemma 2, ensures that $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
@@
Note that $\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cup\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$. For all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$, where $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ with $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})=\Pi(\mathbf{x},\mathcal{O}_{i})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, the control vector (12a) is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. Now, for any $\mathbf{p}\in\mathbb{R}^{3}$ and a unit vector $\mathbf{q}\in\mathbb{S}^{2}$, it can be shown that $\mathbf{p}^{\top}\mathbf{P}(\mathbf{q})\mathbf{p}\geq 0$. Therefore, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\geq 0$. This implies that $\mathbf{u}(\xi)\in\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, and condition (45) holds true.
@@
Next, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, the control vector (12a) is given by $\mathbf{u}(\xi)=-\kappa_{r}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. As mentioned earlier, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\geq 0$. Therefore, $\mathbf{u}(\xi)\in\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, and condition (45) holds true. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$, one can conclude that
$\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})).$ (46)
Next, we show that when the control input corresponds to the obstacle-avoidance mode, $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$. When the control input corresponds to the obstacle-avoidance mode, it is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{a})$, which, as per (13), can be expressed as a linear combination of the vectors $\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$ and $\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. Since $\mathbf{P}(\mathbf{a})$, defined in (14), is an orthogonal projection operator and $\mathbf{0}\in\mathcal{P}(\mathbf{h},\mathbf{a})$, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one has $\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Additionally, one can show that $\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Therefore, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one has $\mathbf{v}(\mathbf{x},\mathbf{a})\in\mathcal{P}(\mathbf{h},\mathbf{a})$. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, using (46), one can conclude that
$\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a}),$
for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$ and claim 1 in Lemma 4 is satisfied.
@@
Next, we proceed to prove claim 2 in Lemma 4 which states that when $\xi(t_{1},j_{1}+1)\in\mathcal{F}_{1}$ for some $(t_{1},j_{1}+1)\in\text{ dom }\xi$, the control input steers the state $\xi$ to the jump set $\mathcal{J}_{1}$ of the obstacle-avoidance mode in finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty.$
@@
Let us define the set $\mathcal{O}_{i}^{\mathcal{S}}=\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, as shown in Fig. 5. Since obstacle $\mathcal{O}_{i}$ is convex, the set $\mathcal{O}_{i}^{\mathcal{S}}$ is also convex. As a result, the target location has a unique closest point on the set $\mathcal{O}_{i}^{\mathcal{S}}$, represented by $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}}).$ Let us define the set $\mathcal{LS}:=\mathcal{L}_{s}(\mathbf{0},\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}}))\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. According to Remark 3, one has $\mathbf{0}\notin\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$. Therefore, it is clear that $\mathcal{LS}\cap\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\neq\emptyset$ for all $\beta\in[0,\gamma]$. Since $\mathbf{0}\in\mathcal{P}(\mathbf{h},\mathbf{a}),$ the line segment $\mathcal{LS}$ belongs to the plane $\mathcal{P}(\mathbf{h},\mathbf{a}).$ Since $\mathcal{LS}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})=\emptyset$, the line segment $\mathcal{LS}$ also belongs to the exit region $\mathcal{R}_{e}$ (18). Since the hit point $\mathbf{h}$ belongs to $\mathcal{R}_{l}^{i}$, the target location $\mathbf{0}$ is closer to the location $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}})$ than to the hit point $\mathbf{h}$. Hence, for a sufficiently small value of $\bar{\epsilon}$, used in (23), one can ensure that the set $\mathcal{LS}$ belongs to the set $\mathcal{J}_{1}^{\mathcal{W}}$ (22).
@@
Now, if one ensures that the state $\mathbf{x}$, which belongs to the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ after time $(t_{1},j_{1})$, in the obstacle-avoidance mode around obstacle $\mathcal{O}_{i}$, eventually intersects the set $\mathcal{LS}$ at some finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$, then it will imply that $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$, and claim 2 in Lemma 4 will be proven. To that end, let us divide the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, as shown in Fig. 5, into 3 separate subsets as follows:
@@
where the sets $\mathcal{S}_{1},\mathcal{S}_{2}$ and $\mathcal{S}_{3}$ are defined as follows:
@@
Figure 5: The partition of the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$.
@@
We show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ belongs either in the set $\mathcal{S}_{1}$ or in the set $\mathcal{S}_{3}$, the control eventually steers the state $\mathbf{x}$ to the set $\mathcal{S}_{2}.$ Then, we show that for all $\mathbf{x}\in\mathcal{S}_{2}$, the control vector $\mathbf{u}(\xi)$ belongs to the open positive half-space $\mathcal{P}_{>}(\mathbf{0},\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi})$. This implies that the state $\mathbf{x}$, which belongs to the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ after time $(t_{1},j_{1})$, in the obstacle-avoidance mode around obstacle $\mathcal{O}_{i}$, is always steered to the open positive half-space $\mathcal{P}_{>}(\mathbf{0},\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi})$ and will eventually reach the set $\mathcal{LS}$ at some finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1).$
@@
First we show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ is either in the set $\mathcal{S}_{1}$ or in the set $\mathcal{S}_{3}$, the control will eventually steer the state $\mathbf{x}$ to the set $\mathcal{S}_{2}.$
When the control input corresponds to the obstacle-avoidance mode and $\mathbf{x}$ belongs to the set $\mathcal{S}_{1}$, the control vector $\mathbf{u}(\xi)$ in (12a) becomes
$\mathbf{u}(\xi)=\kappa_{r}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi},\kappa_{r}>0.$ (49)
Let $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$ for some $\beta\in[0,\gamma_{a}]$. We know that for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$, the tangent cone to the set $\mathcal{N}_{\gamma-\beta}(\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}))$ at $\mathbf{x}$ is given by
$\mathbf{T}_{\mathcal{N}_{\gamma-\beta}(\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}))}(\mathbf{x})=\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi}).$